# Is this chart lying to me? Automating the detection of misleading visualizations

### v1

@@ Appendix F Examples of language and image type diversity in Misviz
Figure 10 shows four instances highlighting the diversity of Misviz. The two visualizations at the top have embedded text written in German and Estonian. The bottom visualizations are photographs of a visualization shown on television screen and on a printed document.
Appendix G Misviz-synth column types
@@
Appendix H Examples of visualizations with axis metadata
@@
Figure 12: Task prompt for zero-shot misleader detection with zero-shot MLLMs.
Appendix I List of random design variations
@@
Bar charts only: sorting bars by values or by category name, hiding the vertical axis, placing labels on top of the bars or within the bars, using horizontal or vertical value labels, variation of bar colors.
@@
Appendix J Zero-shot prompts
Figure 12 shows the task prompt provided to MLLMs to detect misleading visualizations. The prompt starts by providing task instructions, followed by the definitions of the misleaders in alphabetical order. For Misviz-synth, the misleader category discretized continuous variable is removed from the prompt.
Appendix K Linter rules
We explain below the linter rules to detect some of the misleaders in Misviz and Misviz-synth. These rules are implemented in Python and were designed on the validation set of Misviz-synth to maximize precision over recall.
Truncated axis A visualization has a truncated axis if one or more of its sorted vertical axes starts with a numerical value strictly higher than 0.
Dual axis A visualization has a dual axes if it has two vertical axes with different axis tick labels.
Inconsistent tick intervals A visualization has inconsistent tick intervals if the distance between the ticks on one or more of its numerical or temporal axes is not constant.
@@
Image 69.6 $\pm$ 0.9 72.2 $\pm$ 0.2 94.3 $\pm$ 1.3 45.6 $\pm$ 0.4 27.1 $\pm$ 1.9 34.0 $\pm$ 2.3 79.2 $\pm$ 0.2 83.8 $\pm$ 0.4 90.9 $\pm$ 1.0 66.0 $\pm$ 0.7 76.8 $\pm$ 0.1 76.8 $\pm$ 0.1
Image w. axis (DePlot[Uncaptioned image]) 64.3 $\pm$ 3.0 71.4 $\pm$ 0.8 84.4 $\pm$ 4.7 46.7 $\pm$ 1.3 22.1 $\pm$ 1.1 27.7 $\pm$ 1.4 81.6 $\pm$ 0.5 86.6 $\pm$ 0.5 90.4 $\pm$ 1.3 71.7 $\pm$ 0.5 80.4 $\pm$ 0.7 80.4 $\pm$ 0.7
Table 9: Average performance with standard deviations (%) of the fine-tuned classifiers on the test sets of Misviz and Misviz-synth for instances with bar, line, or pie charts.
Appendix L Results with standard deviations of fine-tuned classifiers
Table 9 reports the average results with standard deviations of the fine-tuned classifiers on the subset of Misviz with bar, pie, and line charts and on Misviz-synth.
Figure 14: Error examples on the test set of Misviz-synth. The predictions made by GPT-4.1 are shown in blue while the ground truth is shown in green.
Appendix M Additional implementation details
Training sets To train the DePlot axis extractor and the MLP classification head, we split the Misviz-synth train set into two parts. The larger portion, representing 80% of the data, is used to fine-tune DePlot. The remaining 20% is used to train the MLP classification head. To increase DePlot’s sensitivity to varying tick intervals and step sizes, we add 3,500 additional synthetic visualizations to the axis extraction train set. These are specifically designed to introduce diverse axis spacings.
Fine-tuned classifiers The MLP classification head is trained for up to 300 epochs with early stopping. We use a batch size of 256, a learning rate of 5e-5, and the Adam optimizer. To account for label imbalance, we apply a weighted loss based on the frequency of each misleader. The MLP has a hidden dimension of 1,024.
Appendix N Misviz results per data source
@@
Gemini-1.5-Flash 74.6 27.1 48.1 44.9 19.5 20.5 57.9 18.5 18.5
Gemini-1.5-Pro 83.4 35.5 51.7 81.0 37.6 38.6 64.4 28.4 28.4
@@
Table 10 reports the performance of zero-shot MLLMs across three subsets of the Misviz test set, corresponding to its three sources of misleading visualizations. Performance varies significantly across subsets, but no single source is consistently easier or harder across all models. This highlights the complementary value of each data source to the overall Misviz benchmark. Interestingly, the subset with the lowest EM is often consistent within the same model family, suggesting shared weaknesses in handling visualizations from the same data source.
Appendix O Error examples
Figures 13 and 14 show example of errors made by GPT-4.1 on the test sets of Misviz and Misviz-synth, respectively. One example is shown for each error category. Each example shows the prediction of GPT-4.1 in blue, the correct answer in green, and the error category in gray.

### v2

@@ Appendix F Examples of language and image type diversity in Misviz
Figure 11 shows four instances highlighting the diversity of Misviz. The two visualizations at the top have embedded text written in German and Estonian. The bottom visualizations are photographs of a visualization shown on a television screen and on a printed document.
Appendix G Extending Misviz with semi-ground-truth data
Lo et al. (2022) scraped a large collection of images to construct their taxonomy of visualization misleaders. The data were obtained via diverse web searches for keywords related to misleading visualizations. Only a small subset of this corpus was manually annotated, while the remaining unlabeled portion constitutes a potentially valuable resource for training or evaluating detection models. However, large-scale human annotation is costly and time-consuming, and the raw corpus is inherently noisy. To explore whether this unlabeled data can be exploited more effectively, we investigate the feasibility of automatically assigning semi-ground-truth labels using the best-performing MLLM, Qwen2.5 VL 72B. As a pilot study, we collected 1000 images from the unlabeled portion of the corpus and annotated them using the MLLM.
The MLLM was assigned two tasks: (1) determining whether an image depicts a data visualization, and (2) identifying the presence and type of misleaders when applicable. We manually evaluated the predictions on a subset of 100 images. For task (1), the model achieved an accuracy of 98%, correctly identifying 41 images as visualizations. Among these 41 cases, 20 were correctly classified as misleading or non-misleading, and 17 received the exact correct label, among the 13 available. A substantial portion of the errors can be attributed to the low resolution of many images, which makes it difficult to accurately read value labels. Overall, these results suggest that while current MLLMs show promise for coarse filtering, their ability to reliably annotate unlabeled visualization corpora remains limited.
Appendix H Misviz-synth column types
@@
Appendix I Examples of visualizations with axis metadata
@@
Figure 13: Task prompt for misleader detection with zero-shot MLLMs.
Task prompt for misleader detection with zero-shot MLLMs You are given a chart (dimensions: {img_dim[0]} x {img_dim[1]}) with potential misleading regions: Please analyze the image to detect misleaders and define bounding box coordinates for any misleading regions. ** Let’s think it step by step! ** Here is the list of potential misleaders and their corresponding misleading regions: - discretized continuous variable: a map displays a continuous variable is transformed into a categorical variable by cutting it into discrete categories, thus exaggerating the difference between boundary cases. The misleading region is the legend of the map. - dual axis: there are two independent y-axis, one on the left and one on the right, with different scales. The misleading regions are the two vertical axes. - inappropriate axis range: the axis range is too broad or too narrow. The misleading region is the vertical axis. - inappropriate item order: instances of a variable along an axis are in an unconventional, non-linear or non-chronological order. The misleading region is (parts of) an axis. - inappropriate use of line chart: a line chart is used in inappropriate or unconventional ways, e.g., using a line chart with categorical variables, or encoding the time dimension on the y-axis. The misleading region is one of the axis. - inappropriate use of pie chart: a pie chart does not display data in a part-to-whole relationship, e.g., its shares do not sum to 100%. The misleading region is the labels on the pie slices. - inconsistent binning size: a variable, such as years or ages, is grouped in unevenly sized bins. The misleading region is one of the axis or the legend for a map. - inconsistent tick intervals: the tick values in one axis are not evenly spaced, e.g., the tick value sequence is 10, 20, 40, 45. The misleading region is (parts of) one of the axis. - inverted axis: an axis is displayed in a direction opposite to conventions, e.g., the y-axis displays values increasing from top to bottom or the x-axis displays values increasing from right to left. The misleading region is one of the axis. - misrepresentation: the value labels displayed do not match the size of their visual encodings, e.g., bars may be drawn disproportionate to the corresponding numerical value. The misleading region involves at least two objects (bar, pie slice) for which the size difference is not proportional to the value label difference. - truncated axis: an axis does not start from zero, resulting in a visual exaggeration of changes in the dependent variable with respect to the independent variable. The misleading region is the starting tick of the axis. - 3d: the visualization includes three-dimensional effects. The misleading region is the 3D area of the chart. Then output a JSON file containing coordinates for the potential misleaders and explanations. *** Instructions: - ** Please analyze the image (dimensions: {img_dim[0]} x {img_dim[1]}) to detect any misleading regions.** - **Provide the misleading region coordinates with the name of the corresponding misleader** - Your response format must strictly follow the example JSON format: “‘ [ "coordinates": [[100, 200], [150, 200],[100, 300], [150, 300]],"misleader": "Truncated axis", "coordinates": [[250, 300], [300, 300],[250, 350], [300, 350]], "misleader": "Misrepresentation"] “‘
Figure 14: Task prompt variant for bounding box prediction.
Appendix J List of random design variations
@@
Bar charts only: sorting bars by values or by category name, hiding the vertical axis, placing labels on top of or within the bars, using horizontal or vertical value labels, variation of bar colors.
@@
Appendix K Zero-shot prompts
Figure 13 shows the prompt to detect misleading visualizations. Figure 14 shows the variant used to generate bounding boxes, which is based on the prompt from Chen et al. (2025).
@@
Image 70.1 $\pm$ 0.9 73.7 $\pm$ 0.5 88.3 $\pm$ 1.8 58.7 $\pm$ 1.0 11.1 $\pm$ 0.7 14.9 $\pm$ 0.8 72.0 $\pm$ 0.5 71.9 $\pm$ 0.9 92.2 $\pm$ 1.3 64.7 $\pm$ 1.5 68.0 $\pm$ 1.2 68.0 $\pm$ 1.2
Image w. axis (DePlot[Uncaptioned image]) 72.8 $\pm$ 1.2 74.6 $\pm$ 0.6 92.0 $\pm$ 1.4 60.9 $\pm$ 1.4 12.3 $\pm$ 0.5 17.1 $\pm$ 0.5 72.5 $\pm$ 0.7 72.1 $\pm$ 0.6 92.6 $\pm$ 0.3 65.2 $\pm$ 1.2 69.5 $\pm$ 0.9 69.5 $\pm$ 0.9
Table 8: Average performance with standard deviations (%) of the image-axis classifiers on the test sets of Misviz and Misviz-synth for instances with bar, line, or pie charts.
Appendix L Linter rules
We explain below the linter rules for detecting some misleaders in Misviz and Misviz-synth. These rules are implemented in Python and were designed on the validation set of Misviz-synth to maximize precision over recall.
Truncated axis A visualization has a truncated axis if one or more of its sorted vertical axes start with a numerical value strictly higher than 0.
Dual axis A visualization has dual axes if it has two vertical axes with different axis tick labels.
Inconsistent tick intervals A visualization has inconsistent tick intervals if the distance between the tick labels or relative positions on one or more of its numerical or temporal axes is not constant.
# Chart types Image Image w. axis (DePlot[Uncaptioned image]) Linter
F1 PM F1 PM F1 PM
1 (1,956 instances) 58.2 15.2 60.7 17.4 35.3 7.9
2 (115 instances) 57.2 6.8 59.5 11.5 46.3 6.1
3 (6 instances) 77.5 0.0 38.9 0.0 33.0 0.0
Table 9: Performance of the image-axis classifiers and the linter on Misviz test (%) as a function of the number of chart types.
# Chart types Image Image w. axis (DePlot[Uncaptioned image]) Linter
F1 PM F1 PM F1 PM
1 (1,228 instances) 46.9 13.1 47.9 14.5 16.8 7.7
2 (203 instances) 46.4 24.3 48.1 32.7 17.8 8.9
3 (10 instances) 64.9 46.7 82.5 30.0 16.7 10.0
Table 10: Performance of the image-axis classifiers and the linter on Misviz test (%) as a function of the number of misleader types.
@@
Gemini-2.5-Flash-Lite 86.8 33.1 60.8 52.8 28.0 28.4 74.0 28.0 29.7
@@
Appendix M Results with standard deviations of image-axis classifiers
Table 8 reports the average results with standard deviations of the image-axis classifiers on the subset of Misviz with bar, pie, and line charts and on Misviz-synth.
Figure 16: Error examples on the test set of Misviz-synth. The predictions made by GPT-4.1 are shown in blue, while the ground truth is shown in green.
Model No definitions, no examples Definitions, no examples Definitions with one example
F1 EM F1 EM F1 EM
InternVL3-38B 54.3 29.2 58.4 25.5 58.5 26.6
Qwen2.5-VL-72B 61.2 15.7 62.8 23.2 62.0 22.1
Table 12: Prompt sensitivity analysis on the val set of Misviz (%).
Appendix N Additional implementation details
Training sets We fine-tune DePlot on the Misviz-synth train set. The train-small set is used to train the classification head.
Image-axis classifiers The classification head is trained for up to 300 epochs with early stopping. We use a batch size of 256, a learning rate of 5e-5, and the Adam optimizer. To account for label imbalance, we apply a weighted loss based on the frequency of each misleader. The classification head has one hidden layer with 1.024 units. During training, we keep track of the best model weights. The best weights are updated at each epoch if the F1 score increases on the validation sets of both Misviz and Misviz-synth. While the classifier is not trained on Misviz, using the Misviz validation set ensures we do not overfit to the type of visualizations shown in Misviz-synth.
Appendix O Detailed Misviz results
Tables 9 and 10 report the performance of the image-axis classifiers and the linter on the Misviz test set as a function of the number of chart types and misleader types present in the visualization, respectively. These results provide further insights into the generalization gap between the Misviz-synth training data and the real-world visualizations of Misviz. Visualizations containing multiple chart types are substantially more challenging. This is expected, as multiple chart types increase visual complexity and introduce interactions that are not represented in Misviz-synth. In contrast to chart types, increasing the number of misleaders improves performance for both classifiers and for the linter. Visualizations with more misleaders increase the likelihood that at least one misleader is detected.
Table 11 reports the performance of MLLMs across three subsets of the Misviz test set, corresponding to its three sources of misleading visualizations. Performance varies significantly across subsets, but no single source is consistently easier or harder. This highlights the complementary value of each data source to the overall Misviz benchmark. Interestingly, the subset with the lowest EM is often consistent within the same model family, suggesting shared weaknesses in handling visualizations from the same data source.
Appendix P Prompt sensitivity analysis
We analyze the sensitivity of the two best-performing open-weight MLLMs to different input prompts. We consider three prompts: (1) the prompt only includes the name of the 12 misleader categories, (2) the prompt includes the definitions, (3) the default prompt, which includes the definitions and, in some cases, an example. The results for the Misviz test set, shown in Table 12, indicate a high level of sensitivity for EM and a moderate level for F1. The best prompt depends on the MLLM. Interestingly, InternVL3-38B achieves the highest EM when only the names of the misleaders are provided, without any additional context.
Appendix Q Error examples
Figures 15 and 16 show examples of errors made by GPT-4.1 on the test sets of Misviz and Misviz-synth, respectively. One example is shown for each error category. Each example shows GPT-4.1’s prediction in blue, the correct answer in green, and the error category in gray.